Assuming that the samples come from normal distribution, find the margin of error E given the following:







n = 16 and x̄ = 50 with s = 4.2, 95% confidence

1
Expert's answer
2022-03-27T16:02:04-0400

"E=\frac{s}{\sqrt{n}}t_{n-1,\frac{1+\gamma}{2}}=\frac{4.2}{\sqrt{16}}\cdot 2.1314=2.23797"